OpenStudy (anonymous):

Find an expression for the mean and standard deviation for the following data.Complete the calculations algebraically. 4a+6b a-b a+b 2a-2b

OpenStudy (anonymous):

you have 4 quantaties so to calculate mean, add those all up then divide by 4: 4a+6b + a-b + a+b + 2a-2b =4a+a+a+2a +6b-b+b-2b =8a + 4b so the mean is (8a+4b)/4 = 2a+b for standard deviation, ta